OpenStudy (skittles_for_life6422):

WILL GIVEN MEDAL AND FAN!!!!!!!!!!!!!!!!!!!! ONLY 5 QUESTIONS

OpenStudy (cutestgirlaround):

whatcha going to do with that dessert

OpenStudy (anonymous):

😂😂😂

OpenStudy (skittles_for_life6422):

In a 30 degree - 60 degree - 90 degree, triangle what is the longer leg?

OpenStudy (anonymous):

for the first question: what is the biggest degree? that is the longer leg

OpenStudy (skittles_for_life6422):

answer choices A) \[\sqrt{2} \] times the length of the hypotenuse B)\[\sqrt{3}\] times the length of the hypotenuse C) 2 times the length of the shorter leg D)\[\sqrt{3}\] times the length of the shorter leg

OpenStudy (skittles_for_life6422):

the biggest one is 90

OpenStudy (anonymous):

then thats the answer you just have t aply it to the one of the choices

OpenStudy (anonymous):

to not t

OpenStudy (skittles_for_life6422):

are you saying that 90 is the answer??

OpenStudy (anonymous):

you have to aply it to the other choices

OpenStudy (skittles_for_life6422):

how exactly do you do that

OpenStudy (anonymous):

ok so A says this: √ 2 times the length of the hypotenuse which really means the square root of 2 times the length of the hypotenuse (which you learned before is 90˚)

OpenStudy (skittles_for_life6422):

ok then what

OpenStudy (anonymous):

then you see if that works and if it doesn't then you go on to the next until one question works

OpenStudy (skittles_for_life6422):

ok i think it is the third but im not sure

OpenStudy (skittles_for_life6422):

is it right

OpenStudy (anonymous):

the third one? well it says 2 times the length of the shorter leg the shorter leg is 30˚ do you think this is the right answer? I am only asking you

OpenStudy (skittles_for_life6422):

umm i well im not sure if actually doing this write

OpenStudy (skittles_for_life6422):

right*

OpenStudy (anonymous):

hmmm me either this question is getting harder and harder as i look at it😂

OpenStudy (skittles_for_life6422):

yeah it is @mathmale or @StudyGurl14 do you know how to do this

OpenStudy (anonymous):

hope you get help soon

OpenStudy (skittles_for_life6422):

thx

OpenStudy (skittles_for_life6422):

@cutestgirlaround do you know how to do this

OpenStudy (cutestgirlaround):

no

OpenStudy (skittles_for_life6422):

@phi @dan815 @LegendarySadist @ChantySquirrel1129** do you know how to do this

OpenStudy (rainbowdashrocks):

whats the question

OpenStudy (rainbowdashrocks):

i am pretty sure 90 would be the biggest

OpenStudy (rainbowdashrocks):

if 90 is bigger than 60 and 30 than yes 90 is correct

OpenStudy (skittles_for_life6422):

yes that is correct

OpenStudy (rainbowdashrocks):

of coarse it is i said it lol

OpenStudy (skittles_for_life6422):

@RAINBOWDASHROCKS 90 that is not one of the answer choices

OpenStudy (studygurl14):

The longest leg is the one across from the 60 degree angle (\(\sqrt{3}\)) The hypotenuse is longer, but it's not considered a leg

OpenStudy (skittles_for_life6422):

how did you get that i don't know how to do this

OpenStudy (studygurl14):